OpenStudy (anonymous):

One number is five more than three times another. If their difference is 35, what is the larger number?

OpenStudy (johnweldon1993):

Let our numbers be 'x' and 'y' one number (x) is 5 more than 3 times the other... 5 more = + 5 3 times other number = 3y so x = 3y + 5 now the difference between these 2 numbers = 35....so x - y = 35 we now know that x = 3y + 5 so substitute that into this new equation...and solve for y... x - y = 35 ...becomes 3y + 5 - y = 35 What does 'y' equal...?

OpenStudy (anonymous):

y=15

OpenStudy (johnweldon1993):

Correct....now that we know y = 15.....we can solve for 'x' we know that x - y = 35 since y = 15 we have x - 15 = 35 what does 'x' equal...?

OpenStudy (anonymous):

x=50

OpenStudy (johnweldon1993):

Right....so we have that x = 50 and y = 15 lets check if it's correct....we know that x = 3y + 5 right? well if x = 50 and y = 15 ...this is written as 50 = 3(15) + 5 is this a correct statement...?

OpenStudy (anonymous):

hm yea? thanks so much:)

OpenStudy (johnweldon1993):

50 = 3(15) + 5 50 = 45 + 5 50 = 50 so yes it is a correct statement...so these numbers do indeed work....so the answer to your question of "what is the larger number" is 50... no problem :)